OpenStudy (anonymous):

Solve the problem. A rain gutter is made from sheets of aluminum that are 25 inches wide. The edges are turned up to form right angles. Determine the depth of the gutter that will allow a cross-sectional area of 50 square inches. There are two solutions to this problem. Round to the nearest tenth of an inch. A) 2.5 in. and 10.0 in. B) 2.2 in. and 22.8 in. C) 3.0 in. and 12.0 in. D) 1.8 in. and 18.2 in.

OpenStudy (ujjwal):

perimeter= 2(l+b)-l=25 (you need to subtract one side since one side will be open) Area= \(l\times b\)=50 Find l and b

OpenStudy (calculusfunctions):

The cross-section is a rectangle. If the sheets are 25 in. wide, then after folding, (25 − 2x) inches will remain. x is the height which is the part that was folded. Do you understand thus far?

OpenStudy (anonymous):

i know i have to do -b\[-b+-\sqrt{b^2-ac}/2a\]

OpenStudy (anonymous):

after x(25-2x)=50 -2x^2+25x-50=0

OpenStudy (anonymous):